OpenStudy (anonymous):

If a line passes through the points (2,2) and (4,5), the equation of the line is y-5=3/2(x-__).

OpenStudy (igreen):

It wants you to find the y-intercept I believe.

OpenStudy (anonymous):

So it's similar to the last problem you helped me with?

OpenStudy (igreen):

Yes.

OpenStudy (anonymous):

Thanks c:

OpenStudy (igreen):

But it's a little different, this is in a different form. It's not y = mx + b.

OpenStudy (igreen):

It's in Point Slope Form(y - y1 = m(x - x1).

OpenStudy (igreen):

Well the answer is very obvious. y1 and x1 are to a POINT ON THE LINE. Look at the value of y1, and look at the points they give you, you can easily see what x1 is(the blank).

OpenStudy (igreen):

y - y1 = 3/5(x - x1) This is the equation without any points in it. y - 5 = 3/5(x - x1) This is the equation with the y1 value inputted in it. These are the points they give you: (2, 2) and (4, 5). Notice the y-values for these two points..one of them is the same as the y value inputted into the equation..so what is the x-value of that point?

OpenStudy (igreen):

(4, 5) The y-value is 5, like in the equation, and the x-value is 4! So x1 = 4, which is what you will put in the blank.

OpenStudy (anonymous):

Wow! That's a lot ._. Well thank you! I think I get it!